A financial advisor invests some money in a municipal bond paying 3% annual interest and $5000 more than that amount in a certificate of deposit paying 4% annual interest. To earn $410 per year in interest, how much should he invest at each rate?
Answer by richwmiller(17219)

x+y=5000,
0.03*x+0.04*y=410
x=5000-y
0.03*(5000-y)+0.04*y=410
150-0.03y+0.04*y=410
0.01*y=260
y=26000 at 4%
x=5000-y
x=-21000 at 3%
check
0.03*-21000+0.04*26000=410
-630+1040=410
410=410
ok